OpenStudy (anonymous):

For all positive integers a and b, let a·b defined by a·b=a^b+1/a-1. What is the value of 4·2 ?

OpenStudy (anonymous):

I get 17/3 a·b=a^b+1/a-1 4·2 =4^2+1/4-1 = 17/3